OpenStudy (tvinsant):

Please help will fan and medal: Factor. 4x² - 10x + 6

OpenStudy (tvinsant):

4x^2 - 10x +6

OpenStudy (trojanpoem):

(2x - 3) ( 2x - 2) = 0 x = 3/2 x= 1

OpenStudy (dtan5457):

Borrow payback method. 4x^2-10x+6 multiply a by c x^2-10x+24 (x-6)(x-4) plug back in a (4x-6)(4x-4) divide by 2 (2x-3)(2x-2)<<<factored

OpenStudy (trojanpoem):

Cool :D

OpenStudy (trojanpoem):

I was going to tell him to use -b +- sqrt of b^2 - 4ac / 2a

OpenStudy (tvinsant):

Oh. The answer choices are: a. (2x+3)(2x-2) b. (2x-3)(2x+2) c. 2(2x-3)(x-1) d. 2(2x-3)(x+1)

OpenStudy (tvinsant):

is the answer d?

OpenStudy (trojanpoem):

check my answer and compare

OpenStudy (trojanpoem):

i mean no c

OpenStudy (tvinsant):

oh but it isn't in any of the answer choices?

OpenStudy (tvinsant):

oh so its c?

OpenStudy (trojanpoem):

(2x - 2) = 2(x-1)

OpenStudy (tvinsant):

ok. thank you :)

OpenStudy (trojanpoem):

yw

OpenStudy (tvinsant):

oh... could you help me with another problem? @TrojanPoem

OpenStudy (trojanpoem):

With no doubt

OpenStudy (tvinsant):

Ok :) thanks. Factor. 6x^2-46x-16

OpenStudy (tvinsant):

a. 2(x+4)(3x-4) b. 2(x+1)(3x-8) c. 2(x-8)(3x-1) d. 2(x-8)(3x+1)

OpenStudy (trojanpoem):

a sec

OpenStudy (trojanpoem):

(3x + 1)(2x - 16) = (3x + 1)(x - 8) * 2

OpenStudy (tvinsant):

Thank you so much. Sorry for taking your time. Have a nice day :)

OpenStudy (trojanpoem):

You are welcome. Enjoy.